OpenStudy (anonymous):

how to solve for X? please! 9^2+(x/2)^2=x^2

OpenStudy (anonymous):

does (x/2)^2 = (x/4)?

OpenStudy (turingtest):

what happened to the earlier version of this Q ? you should just bump it. Deleting and reposting is prohibited

OpenStudy (turingtest):

no, you need to square the x as well

OpenStudy (anonymous):

so.. x^2-(x^2/4)=81?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

now multiply both sides by 4 to get rid of the fraction, and combine like terms

OpenStudy (turingtest):

...or at least that's one way to do it :)

OpenStudy (turingtest):

whomever is typing please do \(not\) provide the final answer; it will be deleted.

OpenStudy (anonymous):

why are you such a rule nazi?

OpenStudy (turingtest):

'cuz I'm a mean person :)

OpenStudy (turingtest):

I am a moderator, hence I am here to enforce the code of conduct of this site

OpenStudy (anonymous):

\[x=\sqrt{108}?\]ok so i got.. \[4x ^{2}-x ^{2}=324\] \[x ^{2}=108\]

OpenStudy (turingtest):

excellent :)

OpenStudy (anonymous):

hey TuringTest hmm i'm sorry, as i registered today. i was just trying to help finding solution but ok i will be more careful after today, thanks for warning me :)

OpenStudy (turingtest):

sorry to seem mean, but I am trying to elevate the level of this site plus, as you may have noticed, that last person provided the wrong answer anyway so the better policy is to teach, not answer

OpenStudy (turingtest):

@clark_boy understood, no problem :)

OpenStudy (anonymous):

yes exactly I understand you, you are right. thank you :)